While on a picnic, you and four friends decide to ride the paddle boats. You find that only one paddle boat is available to rent. 3 People can ride at one time. How many different sets of 3 people can be chosen from your group to ride the paddle boat?
I think the answer is 10 by listing all the groups. But I don't know if there is a formula to solve this.

Answer by stanbon(48558)

While on a picnic, you and four friends decide to ride the paddle boats. You find that only one paddle boat is available to rent. 3 People can ride at one time. How many different sets of 3 people can be chosen from your group to ride the paddle boat?
-------------------------
5C3 = 5!/[(5-3)!*3!] = 5*4*3/1*2*3 = 20/2 = 10
-------------------------
5C3 is " the number of sets of 3 in a group of 5"
-------------------------
You have it right.
Cheers,
Stan H.